# Comparing Human-Centered Language Modeling: Is it Better to Model Groups, Individual Traits, or Both?

### v1

@@ 5.2 Error Analysis and Disparity
We perform a set of error analysis by comparing performance metrics of HaRT and GRIT models (pre-trained with 4 training blocks) for the user-level regression tasks of age and openness prediction across different groups based on a demographic factor. The different groups are created by sampling the test set into the following age buckets: below 18, 18-21, 21-30, 30-45, and above 45. Table 4 shows GRITage performs better for the task of estimating age across all age groups i.e., exhibits lesser error. We also see lesser errors in GRIT models for the openness assessment tasks (Appendix Table 6) as well as conforming results on both tasks when comparing using the MSE metric (Appendix Tables 7 and 8).
Additionally, we use the error analysis results to compare the error disparity Shah et al. (2020) in GRIT models and HaRT. Error disparity can be exemplified by the "Wall Street Journal Effect" – a systematic difference in error as a function of demographics Hovy and Søgaard (2015). It can be calculated as the difference in the computed metric across different groups based on a demographic factor Shah et al. (2020). We compute the mean error disparity ($MED$) as the sum of the differences in the metric (Pearson correlation for age, and disattenuated Pearson correlation for openness) computed for each group averaged by the number of difference pairs.
Table 8 reports the $MED$ for each model-task pair for HaRT, GRITage, and GRITope models, and age estimation and openness assessment tasks. We find GRIT models to demonstrate lower mean error disparity for each metric i.e., making less error as a function of the age groups.
Age bucket
#Users
HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
<18
503
0.223
0.394
0.393
18-21
560
0.230
0.278
0.276
21-30
563
0.512
0.531
0.519
30-45
249
0.485
0.530
0.520
45+
68
0.106
0.205
0.180
Table 4: Pearson $r$ for age over five age buckets using different types of human contexts for error analysis. Bold indicates best in row.
@@
Age ($r$)
0.215
0.181
0.185
OPE ($r_{dis}$)
0.075
0.090
0.072
Table 5: Mean error disparity for age over five age buckets. Bold indicates best in column (lower is better).
@@ 6 Related Work
People use language to communicate and convey meaning more than mere words. Much work in human-centered NLP has focused on identifying and evaluting including human context in our models. Initial studies experimented with grouping people by socio-demographic factors like age or gender (Volkova et al., 2013; Hovy, 2015) and geographical region (Bamman et al., 2014; Garimella et al., 2017) to capture the variation in language usage and meaning among different groups. These works improved sentiment analysis, polarity classification, and topic detection. Other researchers explored human factors like social networks (Huang et al., 2014; Radfar et al., 2020), occupation (Preoţiuc-Pietro et al., 2015), personality (Schwartz et al., 2013; Lynn et al., 2017), and social media attributes (Bamman and Smith, 2015; Lynn et al., 2019) to improve toxic language detection, sarcasm detection, or stance detection.
Some studies go beyond explicit groups and learning individual representations latently or via historical language. Jaech and Ostendorf (2018) learned latent user embeddings for search query completion. Delasalles et al. (2019) conditioned a language model on a recurrently updated latent author representation. Welch et al. (2020) motivate personalized word embeddings by jointly learning a latent representation for each user and generic word representations. Hofmann et al. (2021) combined learned latent representations of the social space with time to produce dynamic contextualized word embeddings. King and Cook (2020) created personalized models using the authors’ historical text. Lynn et al. (2020) attend to user’s past messages to better predict user personality.
Research on adapting pre-trained language models to socio-demographic factors has been minimal. Guda et al. (2021) propose EMPATH-BERT, a demographically-aware model to predict empathy and distress better. Lauscher et al. (2022) probe PLMs to understand if their representations encode socio-demographic information. Hung et al. (2023) generalize the task-specific EMPATH-BERT to create a PLM injected with demographic group information using a dynamic multi-task learning setup. We adapt their mono-lingual BERT-based model to age with out-of-domain data for our comparison.
Several studies (Li et al., 2021; Mireshghallah et al., 2022; Zhong et al., 2021) have explored adapting pre-trained Transformer-based language models to individual human contexts for downstream tasks. Li et al. (2021) find benefits in adding user IDs to the input to generate explanations for recommender systems. Zhong et al. (2021) learned a latent user-specific vector prepended to the input embeddings to classify sentiment better, similar to Mireshghallah et al. (2022), who use static user text identifiers instead. Matero et al. (2021) perform masked language modeling on users’ past messages with message-level attention, producing efficient document representations for stance detection. Soni et al. (2022) propose human language modeling, where language is modeled conditioned on a dynamic user state derived from temporally ordered past user utterances. We use their state-of-the-art model HaRT in our comparison and as a base for our pre-trained model with individual and group human context.
@@ 7 Conclusion
NLP benefits from modeling group traits like sociodemographic factors and individual users in terms of latent human context. However, humans exhibit varying degrees of group and individual characteristics. Through a comparative study of five user- and document-level tasks, we uncover how using individual traits and group characteristics in PLMs optimizes user-level regression tasks like age estimation and openness assessment. Meanwhile, individual human context training alone appears to bolster single-document annotated classification tasks like stance and topic detection. Despite our progress, our research reveals there are still considerable strides to be made in modeling human factors in language models. Our findings provide valuable insight into including human context in pre-trained language models to suit specific applications.
@@ A.1 Experimental Settings
Age bucket
#Users
HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
<18
503
0.627
0.644
0.618
18-21
560
0.557
0.608
0.592
21-30
563
0.715
0.741
0.738
30-45
249
0.594
0.669
0.667
45+
68
0.567
0.546
0.599
@@
Age bucket
#Users
HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
<18
1113
4.07
2.52
2.82
18-21
1387
6.52
4.00
3.89
21-30
1557
17.82
12.64
13.11
30-45
695
48.59
39.79
40.43
45+
248
114.92
121.66
134.72
@@
Age bucket
#Users
HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
<18
503
0.423
0.410
0.429
18-21
560
0.496
0.487
0.506
21-30
563
0.429
0.380
0.381
30-45
249
0.578
0.489
0.489
45+
68
0.584
0.501
0.467

### v2

@@ 5.2 Error Analysis and Disparity
We conduct an error analysis as a function of a socio-demographic group attribute. Specifically, we measure GRIT and HaRT for error disparity (Shah et al., 2020) – – a systematic difference in error as a function of demographics as exemplified by the “Wall Street Journal Effect” Hovy and Søgaard (2015). We analyzed outcome and error disparity on age and openness prediction from both the models with individual context (HaRT) and combined individual plus group context (GRIT).
First, we split the task test dataset into different buckets based on the age groups (specifically, <18, 18-21, 21-30, 30-45, and >45 years old) of the users in the test set. Then, we compare the performance of our models across these buckets. Results from table 4 indicate that pre-training with individual and group context together performs better for estimating age across all the age groups, which implies it makes fewer errors as a function of the socio-demographic attribute age. We see similar trends for assessing openness personality (Appendix Tables 6, 8). This suggests that the group attribute prediction maybe acting as a regularizer for models pre-trained with both individual and group contexts, thus aiding the models to make fewer errors across all age buckets.
To further confirm, we compute the mean error disparity ($MED$) as the sum of the differences in the performance metric (Pearson correlation for age, and disattenuated Pearson correlation for openness) across each pair of age buckets, that is averaged by the number of pairs (Shah et al., 2020). A lower averaged sum of differences implies lesser errors as a function of the age groups. Lower $MED$ scores for models pre-trained with individual and group context in Table 5 support our previous error analysis.
Age bucket #Users HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
<18 1113 0.223 0.394 0.393
18-21 1387 0.230 0.278 0.276
21-30 1557 0.512 0.531 0.519
30-45 695 0.485 0.530 0.520
45+ 248 0.106 0.205 0.180
Table 4: Pearson $r$ for age over five age buckets using different types of human contexts for error analysis. Bold indicates best in row. We find no statistical difference between GRITage and GRITope for buckets 21-30 and 30-45. All other results show statistical significance $p<0.05$ using paired t-test.
@@
Age ($r$) 0.215 0.181 0.185
OPE ($r_{dis}$) 0.075 0.090 0.072
Table 5: Mean error disparity for age estimation and openness personality assessment over five age buckets. Bold indicates best in column (lower is better).
@@ 6 Related Work
Much work in human-centered NLP has focused on identifying and evaluating inclusion of human context in our models. Initial studies experimented with grouping people by socio-demographic factors like age or gender (Volkova et al., 2013; Hovy, 2015) and geographical region (Bamman et al., 2014; Garimella et al., 2017) to capture the variation in language usage and meaning among different groups. These works improved sentiment analysis, polarity classification, and topic detection. Other researchers explored adapting to user factors (Lynn et al., 2017), social networks (Huang et al., 2014; Radfar et al., 2020), and social media attributes (Bamman and Smith, 2015) to improve downstream tasks like sarcasm detection, and toxic language detection. Some studies go beyond explicit groups and learning individual representations latently Jaech and Ostendorf (2018); Delasalles et al. (2019) or via historical language Matero et al. (2021).
With respect to pre-trained LMs, recent studies have used adapter-based methodology (Li et al., 2021; Zhong et al., 2021) to include individual human contexts for downstream tasks. More recently, large language models have used user-centric prompting (Li et al., 2023) to include human context. However, few studies have explored including human context within the pre-training regime of LMs. Hung et al. (2023) generalize the task-specific EMPATH-BERT (Guda et al., 2021) to create a PLM injected with demographic group information using a dynamic multi-task learning setup. Further, Soni et al. (2022) pre-train a LM with individual human context derived from user language. Our study aims at comparing the impacts of pre-training LMs with individual, or group, or combined individual plus group human context.
@@ 7 Conclusion
NLP benefits from modeling latent human context, such as sociodemographic group traits or individual tendencies. A recent development has been to incorporate this additional human context into the PLMs’ pre-training regimen. However, humans exhibit varying degrees of group and individual characteristics. Learning about the impacts of pre-training with different types of human context will propel us forward in including human context in our base LLMs (Soni et al., 2023). To assess its impacts, we compare three types of PLMs pre-trained with sociodemographic group attributes, individual human contexts, or a combination across five user- and document-level tasks. We find that pre-training with individual and group human context improves the two user-level tasks: age and personality prediction. Pre-training with individual human context improves the three single-document classification tasks, including stance and topic detection. Pre-training solely on group context helps group-based document classification tasks, though not optimally. Our findings show a promising step towards modeling human context and provide valuable insights for the NLP community to think of additional strategies for improving models with task-dependent human context in pre-training.
@@
Acknowledgments
This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #585968. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ A.1 Experimental Settings
Age bucket #Users HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
<18 503 0.627 0.644 0.618
18-21 560 0.557 0.608 0.592
21-30 563 0.715 0.741 0.738
30-45 249 0.594 0.669 0.667
45+ 68 0.567 0.546 0.599
@@
Age bucket #Users HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
<18 1113 4.07 2.52 2.82
18-21 1387 6.52 4.00 3.89
21-30 1557 17.82 12.64 13.11
30-45 695 48.59 39.79 40.43
45+ 248 114.92 121.66 134.72
@@
Age bucket #Users HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
<18 503 0.423 0.410 0.429
18-21 560 0.496 0.487 0.506
21-30 563 0.429 0.380 0.381
30-45 249 0.578 0.489 0.489
45+ 68 0.584 0.501 0.467